OpenStudy (anonymous):

what are the real or imaginary solutions of the polynomial equation125x^3+343=0

OpenStudy (amistre64):

might want to consider this as a cubic sum formula

OpenStudy (amistre64):

sum of cubics that is

OpenStudy (anonymous):

Take 343 to the LHS, then divide both sides by 125. What we get is x^3 = -343/125

OpenStudy (anonymous):

We take the cube root of the LHS, and replace root -1 as i. (so as to get imaginary root)

OpenStudy (alexandervonhumboldt2):

7^3=343